# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v2

@@ Abstract
Purpose: This paper introduces a novel graph-based method, GARG-AML, for efficient and effective anti-money laundering (AML). It quantifies smurfing risk, a popular money laundering method, by providing each node in the network with a single interpretable score. The proposed method strikes a balance among computational efficiency, detection power and transparency. Different versions of GARG-AML are introduced for undirected and directed networks.
Methodology: GARG-AML constructs the adjacency matrix of a node’s second-order neighbourhood in a specific way. This allows us to use the density of different blocks in the adjacency matrix to express the neighbourhood’s resemblance to a pure smurfing pattern. GARG-AML is extended using a decision tree and gradient-boosting classifier to increase its performance even more. The methods are tested on synthetic and on open-source data against the current state-of-the-art in AML.
Findings: We find that GARG-AML obtains state-of-the-art performance on all datasets. We illustrate that GARG-AML scales well to massive transactions graphs encountered at financial institutions. By leveraging only the adjacency matrix of the second-order neighbourhood and basic network features, this work highlights the potential of fundamental network properties towards advancing fraud detection.
Originality: This paper uses only basic network features and expert knowledge on smurfing to construct a performant AML system. The originality lies in the translation of smurfing detection to these features and network representation. Our proposed method is built around the real business needs of scalability and interpretability. It therefore provides a solution that can be easily implemented at financial institutions or incorporated in existing AML solutions.
@@ 1 Introduction
Money laundering disguises illegally-obtained funds, often from criminal activities, to make them appear legitimate. Once cleaned, the money can be used to finance further crimes. Estimates put the amount of laundered money at $2\%$ to $5\%$ of global GDP, which amounts to USD 2 trillion annually (United Nations Office on Drugs and Crime, (n.d.). Since laundering money involves interactions with and through the normal financial system, regulators have imposed stringent anti-money laundering (AML) rules on financial institutions (FATF, 2012-2023; Levi & Reuter, 2006).
Money laundering involves three main steps (United Nations Office on Drugs and Crime, (n.d.; Levi & Reuter, 2006). Placement refers to the introduction of funds into the financial system. During layering, money is moved around to obscure its illegal origins. Finally, integration is the incorporation of laundered money into broader circulation. Because of the involvement of different steps and actors, network analytics has become an important tool for AML (Deprez et al., 2024).
A popular money laundering method is smurfing (Madinger, 2012; Welling, 1989; Starnini et al., 2021; Zhdanova et al., 2014). Large transactions between two accounts are often flagged according to business rules (Bolton & Hand, 2002; Senator et al., 1995; Oztas et al., 2024) in accordance with regulation (Welling, 1989). Therefore, transactions are split into smaller amounts and transferred using multiple money mules—also called smurfs—making detection of illicit money flows harder since (1) the origin of the funds is obscured and (2) smaller transactions seem less suspicious.
Two types of patterns are typical for smurfing (Starnini et al., 2021): (1) scatter-gather and (2) gather-scatter (Altman et al., 2023). Both patterns are visualised in Figure 1. With scatter-gather, money is sent to its destination account via money mules. Gather-scatter, on the other hand, consists of collecting money from multiple accounts into one (or a few) account(s) and sending it to many destination accounts. Gather-scatter is a popular money laundering method for cryptocurrencies (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021).
Figure 1: The two smurfing patterns described in the literature (Starnini et al., 2021; Altman et al., 2023); as scatter-gather (left) and gather-scatter (right).
To fight money laundering, AML practices use customer and transaction data to estimate each client’s money laundering propensity. When suspicion is raised, investigators build a suspicious activity report (SAR) to file with the regulator (FATF, 2012-2023; Levi & Reuter, 2006). Given the vast volume of monitored transactions and the involvement of different business departments in filing a SAR, two aspects are key to ensuring the effective business adoption of a method:
@@
Despite these business requirements, the current AML literature often focuses on deep graph neural networks (GNNs) (Deprez et al., 2024). GNNs are, however, not interpretable, prohibiting the adoption of these novel network analytics methods in practice. Consequently, many institutions still rely on business rules and manual feature engineering based on expert knowledge, which are expensive to maintain and easy to evade.
Given the specific nature of smurfing and the operational requirements faced by financial institutions, we introduce Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), a novel method that grades clients on the basis of their second-order neighbourhood in the transaction network. GARG-AML operates by analysing the resulting adjacency matrix and computing scores based on the contrast between high-density and low-density regions to quantify the network’s resemblance to smurfing. We introduce two versions of GARG-AMM, one based on directed and one on undirected transaction networks.
Moreover, we present an extension to the base GARG-AML score, increasing smurf detection power by combining GARG-AML with classic network centrality measures and tree-based learners. This shows the versatility and power of our method, without compromising on the business needs of speed and simplicity.
The performance of GARG-AML and its extensions is evaluated against state-of-the-art smurfing detection methods using both synthetic and two open-source network datasets. These datasets, which include millions of clients and transactions, closely resemble real-world financial networks. As a result, our experiments assess both detection accuracy and computational efficiency.
@@
  • We introduce a novel smurfing detection method that is scalable to large transaction networks, interpretable and easily integrated into existing machine learning models;
  • We evaluate our method on multiple datasets which are uniquely suited for assessing the effectiveness of our approach;
  • We compare the current state-of-the-art smurfing detection methods, and we perform the first analysis of the scalability of these methods to very large transaction datasets;
  • Complete implementations of the methods and experiments are provided on GitHub to promote further adoption of and experimentation on the GARG-AML scores by peer researchers and practitioners.
The remainder of this work is organised as follows. Related work on anti-money laundering and smurfing is discussed in Section 2. The methodology behind the GARG-AML score calculation is described in Section 3. Section 4 outlines the experimental evaluation, and the results are presented in Section 5. We discuss the main limitations, in terms of both the data in general and GARG-AML in particular, in Section 6. Section 7 concludes this work.
@@ 2 Related Work
Anti-money laundering (AML), like other areas of financial fraud detection, traditionally relied on IF-THEN business rules—combinations of expert knowledge and regulatory requirements (Bolton & Hand, 2002; Leonard, 1995; Sánchez et al., 2009; Welling, 1989; Cardoso et al., 2022; Deprez et al., 2024). These rules typically monitor individual transactions or aggregated amounts over time (FATF, 2012-2023). While offering transparency in decision-making (Zhu & Hu, 2013), they are costly to maintain (Van Belle et al., 2023) and can be easily circumvented once exposed. As a result, machine learning has emerged as a preferred approach, offering scalability and the ability to detect complex laundering patterns (Chen et al., 2018).
@@
To capture interactions among people, network analytics for AML has been studied extensively in recent years (Deprez et al., 2024). Research has often relied on manual network feature engineering for profile definition. Dreżewski et al. (2015) used different centrality measures, which resulted in nine roles in a criminal network. Fronzetti Colladon & Remondi (2017) used similar metrics across industries and analysed the correlation with the target label to demonstrate the added value of network analytics for AML. Wu, Liu, Chen, Huang, Zheng & Zhang (2021) define temporal motives specifically tailored to money laundering and mixing detection in the Bitcoin network.
@@
In addition to manual feature engineering, graph neural networks (GNNs) have also been studied extensively (Lo et al., 2023; Alarab et al., 2020; Weber et al., 2019; Deprez et al., 2024). These methods capture the complex geometry of suspicious money laundering activities in financial networks by learning network topology while also incorporating node features. A drawback is that GNNs are not interpretable (Kute et al., 2021), which limits their use by investigators when filing their reports.
@@
Smurfing has long been a popular laundering method (Welling, 1989; Senator et al., 1995). Welling (1989) suggested that it has been applied since 1970, when the Bank Secrecy Act was passed in the USA, which required reporting of transactions above $\mathdollar 10,000$. To this day, similar rule-based systems are in place that require immediate investigation of transactions that surpass a given amount (Senator et al., 1995; Oztas et al., 2024; Bolton & Hand, 2002).
Research has already been conducted on smurfing detection. Zhdanova et al. (2014) introduced PSA@R for detecting smurfing in mobile money networks. It uses different underlying workflow models. Transaction streams that differ from the process model are flagged as anomalous. Additional criteria are applied to find money mules on the basis of the difference in the amounts of money coming into and going out of an account.
FlowScope (Li et al., 2020) tries to find anomalous sub-graphs on the basis of money flow. The authors constructed a k-partite network with two sets of nodes: nodes with net inflow to or net outflow from the bank and the remaining nodes, which correspond to the middle men. Li et al. (2020) defined their anomaly score on the basis of the difference between the inflow and outflow of money at these middle men, and they claimed that smurfs transfer almost all the money they receive.
A similar idea was put forward with AutoAudit (Lee et al., 2020). The full adjacency matrix of the network is reordered, and depending on the density of the different blocks, a group of transactions is flagged as smurfing.
Starnini et al. (2021) presented a method that merges two transaction datasets and retains only those connections where both transactions occur within a specific time window. Prior to this merger, a substantial amount of data, i.e., nodes deemed trustworthy by the bank together with a majority of transactions, is discarded. After merging, the method retains certain motifs and applies an additional post-processing step, which then triggers manual investigations by AML experts.
SMoTeF (Shadrooh & Nørvåg, 2024) is a three-step approach. The first is smurfing pattern extraction, using existing methods. The novelty comes from the second and third steps, where the patterns are checked against temporal constraints and maximal money flow, respectively. Many transactions are again pruned before making final predictions.
@@
Closely related to smurfing is mixing, which is a money laundering method that is used for cryptocurrencies (Liu et al., 2021). Coins are transferred to the same entity and are redistributed afterward (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021). This obfuscates the history of ownership, which results in an inability to know the money’s real origin. Mixing can be viewed as the second smurfing pattern in Figure 1, i.e., gather-scatter.
An anomaly score for mixing detection was introduced by Prado-Romero et al. (2018). The authors claimed that mixing services act as bridges between communities in a transaction network by linking otherwise unrelated users. After community detection, the anomaly score expresses the difference in a node’s intercommunity links compared with those of the other nodes in the same community.
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) constructed two networks for network feature extraction. The first links accounts using their transactions, whereas the second network represents both accounts and transactions as nodes to leverage transaction information. The extracted network features are based on predefined temporal motifs, which are added to the account and transaction features to detect smurfing in a positive and unlabelled (PU)-learning setting.
The above literature is summarised and juxtaposed against our proposed method according to the methodology in Table 1 and according to the experimental set-up in Table 2. As expected, most methods for anti-money laundering on fiat currency consider only scatter-gather patterns, while research on crypto-currency is mostly concerned with gather-scatter.
The evaluation in most papers is done on relatively small datasets, making it harder to judge whether they scale well to the massive transaction networks faced by financial institutions. The dataset that was used in most of the above papers is the Czech dataset (Data World, 1999), which is a rather small network that contains approximately 11,000 nodes and approximately 273,000 edges. This dataset is unlabelled, so the authors of each paper inject synthetic pattern. Additionally, in most papers, evaluation and comparison is limited to threshold-dependent metrics like precision, recall and F1.
@@
Shadrooh & Nørvåg (2024) Scatter-gather
@@
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) Gather-scatter
@@
Shadrooh & Nørvåg (2024) Czech Financial & Enron & Darpa 11,374 & 16,771 & 371,000 Time, Precision, Recall, F1
@@
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) WalletExplorer $\pm 2,500,000$ TPR, FPR, geometric mean
@@
In this paper, we propose a new method that scores all clients individually, where score can be calculated in parallel. These score can also be used both for scatter-gather as well as for gather-scatter patterns. The experiments are done on very large, realistic data, where we use threshold-independent metrics that allow us to have a comprehensive comparison among the methods under consideration.
@@ 3 GARG-AML
The main idea of GARG-AML is to provide each individual node with a score expressing how likely it is that this entity is involved in a smurfing scheme. This is done by analysing key characteristics of the second-order neighbourhood of the nodes. Below, we start by formally defining what we mean with second-order neighbourhood. Afterwards, we provide two version for GARG-AML, one for undirected and one for directed networks. The full GARG-AML method is given in Algorithm 1 in the experimental set-up in Section 4.2.
@@ 3.1 Preliminaries and Notation
A network $G(V,E)$ is defined by a set of nodes, $V=\{v_{1},\ldots,v_{k}\}$, and a set of edges, $E\subset V\times V$. The edges can be represented using the adjacency matrix, $A$, where the elements are defined as
@@
Suppose $\mathcal{N}_{k}(v)$ represents the set of nodes that are at distance exactly $k$ from node $v$, with $\mathcal{N}_{0}(v)=\{v\}$. For GARG-AML, we consider the neighbours of each node $v$ at distances one and two, i.e., $\mathcal{N}_{1}(v)$ and $\mathcal{N}_{2}(v)$, respectively. Suppose that $\mid\mathcal{N}_{1}(v)\mid=n$ and $\mid\mathcal{N}_{2}(v)\mid=m-1$; then, the adjacency matrix of the second-order neighbourhood of $v$ is of size $(m+n)\times(m+n)$.
@@
3.2 Undirected
We introduce GARG-AML using the smurfing pattern defined by Starnini et al. (2021) and illustrated in Figure 2. In this example, we want to give node $A$ a score. The adjacency matrix of its second-order neighbourhood, given in Equation (1), is constructed by arranging the nodes in a specific order: node $A$ is placed first, followed by the second-order neighbour $E$, and then its first-order neighbours $B,C$ and $D$. This ordering results in an adjacency matrix that is divided into four distinct blocks. The diagonal blocks contain only zeros, whereas the off-diagonal blocks contain ones.
In this example, we will consider only the scatter-gather pattern, as this is most widely recognised as smurfing. The GARG-AML calculations and node order stay the same in the gather-scatter case.
Figure 2: Example of a smurfing pattern. AA is the node that will receive the score. The first-order neighbours ($B,C,D$) are shown in purple, and the second-order neighbour ($E$) is shown in green.

### v3

@@ Abstract
Purpose: We introduce GARG-AML, a fast and transparent graph-based method to catch ‘smurfing’, a common money-laundering tactic. It assigns a single, easy-to-understand risk score to every account in both directed and undirected networks. Unlike overly complex models, it balances detection power with the speed and clarity that investigators require.
Methodology: The method maps an account’s immediate and secondary connections (its second-order neighbourhood) into an adjacency matrix. By measuring the density of specific blocks within this matrix, GARG-AML flags patterns that mimic smurfing behaviour. We further boost the model’s performance using decision trees and gradient-boosting classifiers, testing the results against current state-of-the-art on both synthetic and open-source data.
Findings: GARG-AML matches or beats state-of-the-art performance across all tested datasets. Crucially, it easily processes the massive transaction graphs typical of large financial institutions. By leveraging only the adjacency matrix of the second-order neighbourhood and basic network features, this work highlights the potential of fundamental network properties towards advancing fraud detection.
Originality: The originality lies in the translation of human expert knowledge of smurfing directly into a simple network representation, rather than relying on uninterpretable deep learning. Because GARG-AML is built expressly for the real-world business demands of scalability and interpretability, banks can easily incorporate it in their existing AML solutions.
@@ 1 Introduction
Criminals launder an estimated USD 2 trillion a year—up to $5\%$ of global GDP—to disguise their illicit profits and fund future crimes (United Nations Office on Drugs and Crime, (n.d.)). To mitigate this, regulators have mandated that financial institutions serve as the primary front in anti-money laundering (AML) efforts (FATF, 2012; Levi and Reuter, 2006).
Financial institutions have put in place AML detection and reporting processes. Alert generation is an automatic process for spotting unusual behaviour, and results in many false positives. Manual investigations filters out legitimate transactions, based on both internal (e.g., transactions and client records) and external data (e.g., public domain information searches on company structure). The remaining alerts are further investigated to build cases, which contain multiple alerts. Expert investigators decide if a suspicious activity report (SAR) is filed to the regulator (Halford et al., 2025; Jullum et al., 2020; FATF, 2012; Levi and Reuter, 2006).
A large portion of alerts come from business rules flagging large transactions amounts in accordance with regulation (Welling, 1989; Bolton and Hand, 2002; Senator et al., 1995; Oztas et al., 2024; Jullum et al., 2020). These rules also consider aggregation attempts, where criminals make many transactions each staying under the reporting threshold. However, the efficacy of these rules is often compromised by the sheer volume and complexity of global transaction data.
A popular method to circumvent these business rules is smurfing (Madinger, 2012; Welling, 1989; Starnini et al., 2021; Zhdanova et al., 2014). Criminals split transactions into smaller amounts and transfer them through money mules—called smurfs—to dodge the rules. Smurfing makes AML harder since (1) the chain between the criminal activity and the criminals’ bank accounts is interrupted, and (2) smaller transactions through distinct accounts seem less suspicious.
Banks monitor millions of transactions and require multiple departments to report suspicious activity. To be useful, any new tool must meet two criteria:
@@
Despite these business requirements, the current network-based AML literature focuses on black box graph neural networks (GNNs) (Deprez et al., 2024). Consequently, many institutions rely on business rules and manual feature engineering based on expert knowledge.
To address the specific nature of smurfing and the operational requirements, we introduce Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), a novel method that grades clients based on their second-order neighbourhood in the transaction network. GARG-AML quantifies how closely an account’s network structure resembles known smurfing patterns.
Moreover, we present an extension using classic network centrality measures and tree-based learners to increase smurf detection power. This shows the versatility and power of our method, without compromising on the business need of speed and simplicity.
The performance is evaluated against state-of-the-art smurfing detection methods using both synthetic and two open-source network datasets. These datasets, which include millions of clients and transactions, closely resemble real-world financial networks. As a result, our experiments assess both detection power and computational efficiency.
@@
  • We introduce a novel smurfing detection method that is scalable to large transaction networks, interpretable and easily integrated into existing models;
  • We evaluate our method on multiple datasets that are uniquely suited for assessing the effectiveness of our approach;
  • We compare the current state-of-the-art in smurfing detection methods, and we perform the first analysis of the scalability of these methods to very large transaction datasets.
The remainder of this work is organised as follows. Related work on anti-money laundering and smurfing is discussed in Section 2. The methodology behind GARG-AML is described in Section 3. Section 4 outlines the experimental evaluation, with results presented in Section 5. We discuss the main data and methodological limitations in Section 6. Section 7 concludes this work.
Complete implementations of the methods and experiments are provided on GitHub [^1] https://github.com/B-Deprez/GARG-AML to promote further adoption of and experimentation on the GARG-AML scores by peer researchers and practitioners.
@@ 2 Related Work
According to Routine Activity Theory (RAT), criminal acts occur when three elements converge: a motivated offender, a suitable target, and the absence of a capable guardian. In the context of money laundering (Cohen and Felson, 1979):
  • Motivated Offenders: Criminal organizations seeking to integrate ”dirty” money into the legitimate economy without triggering regulatory alarms.
  • Suitable Targets: Financial institutions, whose systemic gaps or high transaction volumes can be exploited to hide illicit flows.
  • Capable Guardians: The AML systems and human investigators tasked with identifying and reporting suspicious activity.
@@
Current technological “guardians” largely rely on IF-THEN business rules that monitor individual transactions or aggregate amounts over time (Bolton and Hand, 2002; Leonard, 1995; Sánchez et al., 2009; Welling, 1989; Cardoso et al., 2022; Deprez et al., 2024). While transparent, these rules are often ineffective as they are easily circumvented once the underlying logic is exposed, reducing their utility as capable guardians (Zhu and Hu, 2013; Van Belle et al., 2023). Additionally, The volume and complexity of monitored transactions often compromise the effectiveness of these guardians (Mat Isa et al., 2025; Lokanan, 2025). As a result, machine learning has emerged as a preferred approach, offering scalability and the ability to detect complex laundering patterns (Chen et al., 2018; Jullum et al., 2020; Ramadhan, 2024).
Network analytics has emerged as a critical tool for AML, shifting the focus from individual transactions to the structural interactions between entities (Deprez et al., 2024). Current literature can be categorized into three primary research streams: manual feature engineering, deep learning architectures, and pattern-specific detection frameworks.
@@
Early research emphasized the use of centrality measures for profile definition. Dreżewski et al. (2015) used centrality measures to define nine roles in criminal networks. Fronzetti Colladon and Remondi (2017) used similar metrics across industries and analysed the correlation with the target label to demonstrate the added value of network analytics for AML. Wu et al. (2021a) defined temporal motives tailored to money laundering and mixing detection in the Bitcoin network.
@@
Deep learning has been studied extensively through GNNs (Lo et al., 2023; Alarab et al., 2020; Weber et al., 2019; Deprez et al., 2024). These methods capture the complex geometry of suspicious money laundering activities in financial networks by learning network topology while also incorporating node features. A drawback is that GNNs are not interpretable (Kute et al., 2021), which limits their use by investigators.
@@
Smurfing has long been a popular laundering method (Welling, 1989; Senator et al., 1995). Welling (1989) sees it as a consequence of the Bank Secrecy Act in the USA, which required reporting of transactions above $\mathdollar 10,000$. Similar monetary-based systems are still in place that require immediate investigation of transactions that surpass a given amount (Senator et al., 1995; Oztas et al., 2024; Bolton and Hand, 2002).
Although many money laundering methods and patterns exist (Suzumura and Kanezashi, 2021), two types of patterns are typical for smurfing (Starnini et al., 2021; Altman et al., 2023): (1) scatter-gather and (2) gather-scatter. Both patterns are visualised in Figure 1. With scatter-gather, money is sent to its destination account via money mules. Gather-scatter, on the other hand, consists of collecting money from multiple accounts into one (or a few) account(s) and sending it to many destination accounts. Gather-scatter is a popular money laundering method for cryptocurrencies (Wu et al., 2021b).
Figure 1: The two smurfing patterns; as scatter-gather (left) and gather-scatter (right).
Smurfing makes use of smurfs, who function as money mules. Smurfs are recruited from outside the criminal organisation for a short period of time. Criminals lure the young or financially vulnerable with promises of quick cash, using them to open accounts in the victim’s name to funnel illicit funds (Esoimeme, 2020; Leukfeldt and Jansen, 2015; Bekkers et al., 2025).
The evolution of laundering techniques like smurfing (or microstructuring) is best explained by Rational Choice Theory (McCarthy and Chaudhary, 2014). Offenders seek to maximize reward while minimizing the risk of detection. By splitting large, reportable sums into smaller transactions funnelled through smurfs, criminals exploit the ‘blind spots’ of rule-based systems. This tactic interrupts the visible chain of funds, making illicit flows appear indistinguishable from legitimate retail banking activity.
@@
Closely related to smurfing is mixing, which is a money laundering method that is used for cryptocurrencies (Liu et al., 2021). Coins are transferred to the same entity and are redistributed afterward, resulting in a gather-scatter pattern.
Prado-Romero et al. (2018) introduce an anomaly score for mixing detection by identifying mixing services that act as bridges between unrelated user communities. Wu et al. (2021a) applies positive and unlabelled (PU)-learning using predefined temporal motifs. These are extracted from two networks, one with just accounts as nodes and the other with accounts and transactions as nodes.
The methodologies and experimental setups of these existing works are summarized and juxtaposed against GARG-AML in Table 1 and Table 2, respectively. Methods for anti-money laundering on fiat currency consider scatter-gather patterns, while research on crypto-currency is mostly concerned with gather-scatter.
The evaluation in most papers is done on relatively small datasets. The Czech dataset (Data World, 1999) is used most. It is a small network containing approximately 11,000 nodes and 273,000 edges. This dataset is unlabelled, so the authors of each paper inject synthetic pattern. Additionally evaluation and comparison is often limited to threshold-dependent metrics like precision, recall and F1. GARG-AML addresses these limitations by providing an interpretable, parallelizable scoring framework evaluated on massive, realistic transaction networks using threshold-independent metric.
@@
Shadrooh and Nørvåg (2024) Scatter-gather
@@
Wu et al. (2021a) Gather-scatter
@@
Shadrooh and Nørvåg (2024) Czech Financial & Enron & Darpa 11,374 & 16,771 & 371,000 Time, Precision, Recall, F1
@@
Wu et al. (2021a) WalletExplorer $\pm 2,500,000$ TPR, FPR, geometric mean
@@ 3 GARG-AML
GARG-AML provides each individual node with a score expressing how likely it is that this entity is involved in a smurfing scheme. This is done by analysing the second-order neighbourhood of the nodes. We start by formally defining some crucial concepts. Afterwards, we introduce two version for GARG-AML, one for undirected and one for directed networks. The full GARG-AML method is given in Algorithm 1.
@@ 3.1 Preliminaries and Notation
During the placement and layering stage, money mules are often used (Raza et al., 2020). These can either be people actively recruited from outside the organisation (Esoimeme, 2020; Leukfeldt and Jansen, 2015) or unaware victims whose bank accounts have been hacked (Raza et al., 2020). Their accounts are used to receive money from criminal activity and to quickly transfer it to the criminal organisation’s account. The use of accounts in the name of outsiders results in the circumvention of strict KYC rules and to put distance between the criminal activity and the criminals’ accounts (Esoimeme, 2020).
One subcategory of money mules is smurfs (Starnini et al., 2021; Zhdanova et al., 2014). Smurf are used to break up large transactions and funnel the money through many accounts in smaller instances. This helps the criminals to avoid business rules that are often only monitoring transactions per counterparty (e.g., between person A and B, and A and C, separately, but not aggregated among A,B and C).
A network $G(V,E)$ is defined by a set of nodes, $V=\{v_{1},\ldots,v_{k}\}$, and edges, $E\subset V\times V$. We represent the network using the adjacency matrix $A$. This elements of $A$ indicate whether pairs of nodes are connected via an edge. Mathematically, it is defined as:
@@
$\mathcal{N}_{k}(v)$ represents the set of nodes that are at distance exactly $k$ from node $v$, with $\mathcal{N}_{0}(v)=\{v\}$. GARG-AML relies on the second-order neighbourhood of node $v$. This is the sub-graph consisting of all nodes at a distance of at most two from $v$, capturing both direct counterparties ($\mathcal{N}_{1}(v)$) and the ‘counterparties of counterparties’ ($\mathcal{N}_{2}(v)$).
Suppose that $\mid\mathcal{N}_{1}(v)\mid=n$ and $\mid\mathcal{N}_{2}(v)\mid=m-1$; then, the adjacency matrix of the second-order neighbourhood of $v$ is of size $(m+n)\times(m+n)$.
@@
3.2 Undirected Network
We introduce GARG-AML using an example based on the smurfing pattern illustrated in Figure 2. We will consider only the scatter-gather pattern, as this is most widely recognised as smurfing. The methodology is the same for gather-scatter patterns.
In this example, node $A$ is given a score. The adjacency matrix of its second-order neighbourhood, given in Equation (1), is constructed by arranging the nodes in a specific order: node $A$ is placed first, followed by the second-order neighbour $E$, and then its first-order neighbours $B,C$ and $D$. This ordering results in an adjacency matrix that is divided into four distinct blocks. The diagonal blocks contain only zeros, whereas the off-diagonal blocks contain ones.
Figure 2: Pure smurfing pattern. AA is the node that is scored. The first-order neighbours ($B,C,D$) are shown in purple, the second-order neighbour ($E$) is shown in green.